# On the Curvature of Regge Metrics

### v1

@@ 2.1 Conditions on Compatible Frames
An orthonormal frame $f=\bigsqcup_{T\subseteq M}f^{T}$ will be compatible if it has some desirable properties that make $\left\langle\left\langle f^{*}\Omega_{\mathrm{dist}},\phi\right\rangle\right\rangle$ both correct from a geometrical standpoint and practical from a computational standpoint. The conditions that a compatible frame must satisfy are fairly technical and may seem arbitrary, so before stating them, we will first provide some motivation.
First, we want the individual vector fields in our frame to be “parallel” across codimension-1 polytopes, in some sense. For each interior codimension-1 polytope $e\subset M$, we can define a frame $E_{e}$ which is orthonormal in the metric $i_{e}^{*}g$; this metric is well-defined since $g$ has single-valued tangential-tangential components on $e$. If $e=T\cap T^{\prime}$, then $E_{e}$ can be extended to two orthonormal frames $E_{e}^{T}$ and $E_{e}^{T^{\prime}}$ by appending outward-facing normal vectors to $e$ for $g|_{T}$ and $g|_{T^{\prime}}$ respectively, which we denote by $\vec{n}$ and $\vec{n}^{\prime}$. Assume that for each $T$, $f^{T}$ can be continuously extended to $\mathring{e}$ by the $C^{1}$ section $f^{T}|_{\mathring{e}}:\mathring{e}\to\mathcal{F}_{O}(T)|_{\mathring{e}}$, and let $\mu_{e}^{T}:\mathring{e}\to O(n)$ be a map such that
@@
While somewhat arbitrary, this notion of “parallelism” is supported by the fact that a piecewise-smooth geodesic, defined as a locally length-minimizing curve, must have a velocity vector that satisfies the same condition we have placed on the $f_{i}$’s.
@@
$=\int_{\mathring{T_{i}}}\left(\left\langle{f^{T_{i}}}^{*}\omega\wedge d\phi\right\rangle+\frac{1}{2}\left\langle{f^{T_{i}}}^{*}[\omega,\omega]\wedge\phi\right\rangle\right)$
@@
As the Riemann curvature should be tensorial, it should always be true that changing the frame by $h$ results in an adjoint action by $h^{-1}$ on $f^{*}\Omega_{\mathrm{dist}}$. For this to be true for all $\phi$, it must be the case that $h^{-1}dh]\!]=0$ along every codimension-1 face $e=T\cap T^{\prime}$, so on $e$, $h^{T}=C_{e}h^{T^{\prime}}$ for some constant matrix $C_{e}$. We will restrict ourselves to the case $C_{e}=I$, since otherwise it would not be possible to modify the transformation $h$ so that it is the identity outside of a small neighborhood of $x_{0}$ but retains the jump condition, losing locality. What this means is that, if any particular frame $f$ is asserted to be compatible, then it is reasonable to assert that in a small enough neighborhood $U$ of a point $x\in\mathring{e}$, the set of compatible frames on $U$ (meaning restrictions of compatible frames to $U$) must be contained in $\{(f|_{U}\cdot h):h\;\text{is piecewise-smooth and continuous}\}$. Therefore, piecewise-smooth frames are usually not compatible, and the set of piecewise-smooth compatible frames must all have the same jump conditions along codimension-1 polytopes—the matrix $\mu_{e}^{T}(\mu_{e}^{T^{\prime}})^{-1}$ cannot depend on $f$. The choice $\mu_{e}^{T}(\mu_{e}^{T^{\prime}})^{-1}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}$ is the simplest, and it is consistent with the case of continuous metrics and frames.
@@
Note that since the final entries of $E_{e}^{T}$ and $E_{p,e}^{T}$ are associated to normal vectors to $e$ in the metric $g_{T}$, $A_{p,e}^{T}=\begin{bmatrix}A_{p,e}&0\\ 0&\pm 1\end{bmatrix}$, where $A_{p,e}$ does not depend on the polytope $T$. Because $E_{e}^{T^{\prime}}$ has a normal vector that points in a different direction to that of $E_{e}^{T}$, but $E_{p,e}^{T^{\prime}}$ does not, we also have $A_{p,e}^{T}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}A_{p,e}^{T^{\prime}}=A_{p,e}^{T^{\prime}}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}$, and therefore
@@
which implies $\sum_{i=1}^{k}\theta_{p}^{T_{i}}=2m\pi$ for some integer $m$. However, in general, this sum can take any positive value if the metric is discontinuous at $p$. Therefore the frame $f^{T}$ cannot always be continuous at $p$. This is the origin of angle defect. In order to control the discontinuity of $f^{T}$ as much as possible, we will restrict it to only rotate at a constant speed, and only in the plane orthogonal to $p$.
To ensure that we are still able to apply integration by parts, despite the fact that the frame is discontinuous at $p$, we will follow the strategy outlined in [1, p. 2] and require that $f^{T}$ has some smoothness and continuity when pulled back to the blow-up $B_{T}$ of $T$. The blow-up $B_{T}$ of $T$ is essentially a polytope that has one codimension-1 face for each codimension-$d$ face of $T$ with $d\geq 1$, with tangencies related to inclusion relations between the original faces. There is a corresponding blow-down map $\Phi^{T}:B_{T}\to T$ which restricts to a diffeomorphism $\Phi^{T}|_{\mathring{B}_{T}}:\mathring{B}_{T}\to\mathring{T}$. The blow-up has an exceptional set which we will call $E_{B_{T}}$, consisting of the faces of $B_{T}$ of codimension $\geq 2$, and it can be safely ignored for the purpose of Riemann integration. Some information on blow-ups of manifolds can be found in the appendix.
@@
A frame $f=\bigsqcup_{T\subseteq M}f^{T}$, where $f^{T}:\mathring{T}\to\mathcal{F}_{O}(T)|_{\mathring{T}}$ is an orthonormal frame for the metric $g^{T}$, is compatible if:
@@
    For each $e\subset T\subseteq M$, let $\mu_{e}^{T}:e\to O(n)$ be the matrix such that $E_{e}^{T}\cdot\mu_{e}^{T}=f^{T}|_{e}$, and suppose $e=T\cap T^{\prime}$. Then $\mu_{e}^{T^{\prime}}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\mu_{e}^{T}$.
@@
    For each $p\subset T\subseteq M$, there exists an orientation-preserving diffeomorphism
    $\psi^{T}_{p}:[0,1]\times p\to\overline{{\Phi^{T}}^{-1}(\mathring{p})}$ such that $\Phi^{T}\circ\psi^{T}_{p}(s,x)=x$. For any $e$ such that $p\subset e\subset T$, there exists a continuous function $r_{p,e}^{T}:p\to\mathbb{R}$ such that
@@
    for all $x\in p$ and all $s\in[0,1]$. In other words, the multi-valuedness of $f^{T}$ at codimension-2 faces is controlled so that $f^{T}$ is continuous when expressed in cylindrical coordinates around $x\in p$ and rotates in the plane orthogonal to $p$ at a rate depending only on $x$. There is no restriction on behavior near faces of higher codimension. Note that $r_{p,e}^{T}=-r_{p,e^{\prime}}^{T}$ if $p=e\cap e^{\prime}$.
@@
As we mentioned above, condition 1 makes it possible to integrate by parts even though the frame may have discontinuities on polytopes of codimension 2 or greater. The fact that $F^{T}$ is Lipschitz and $C^{2}$ on the set $B_{T}\backslash E_{B_{T}}$ is necessary for the identity $d{F^{T}}^{*}\omega={F^{T}}^{*}d\omega$ and for ${F^{T}}^{*}\omega$ to be bounded and continuously extendable to $B_{T}\backslash E_{B_{T}}$. These are all necessary conditions for Stokes’ Theorem to hold for the form ${F^{T}}^{*}(\omega\wedge\pi^{*}\phi)$ on $B_{T}$. Conditions 2 and 3 are necessary for the frame $f$ to have some semblance of continuity and for the distributional curvature to transform like a tensor, as mentioned previously, and condition 4 states that compatible frames can be “smoothed out” without introducing singularities.
@@
Having said all of this, one still might wonder if the use of $B_{T}$ could be avoided by performing integration by parts on $\overline{f^{T}(\mathring{T})}$ instead. After all, we are allowing $f^{T}$ to have discontinuities on codimension-2 faces of $T$, so it is reasonable to expect the boundary of $\overline{f^{T}(\mathring{T})}$ to have “extra” codimension-1 faces much like $B_{T}$ does. Unfortunately $\overline{f^{T}(\mathring{T})}$ may fail to be a Whitney manifold. For instance, this could happen if $r_{p,e}^{T}$ vanishes on a disconnected set which has nonzero measure in $p$. In a situation like this, integration by parts on $\overline{f^{T}(\mathring{T})}$ is infeasible. However, by pulling back forms by $F^{T}$ to the manifold $B_{T}$, which is regular enough to permit an integration by parts formula, we regain this ability.
@@ 2.2 The Integration by Parts Step
For $d>2$, the integrals over $\Phi^{-1}(\mathring{\Delta}_{d})$ vanish for any codimension-$d$ face $\mathring{\Delta}_{d}\subset S_{d}(T)$.
Indeed, $\phi^{j}_{i}$ is an (n2)(n-2)-form, so its trace $i_{\mathring{\Delta_{d}}}^{*}\phi^{j}_{i}$ vanishes on any $\mathring{\Delta}_{d}$ with $d>2$. Since $i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}F^{*}\pi^{*}=i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}\Phi^{*}=\Phi^{*}i_{\mathring{\Delta}_{d}}^{*}$, we have $i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}F^{*}\pi^{*}\phi^{i}_{j}=\Phi^{*}i_{\mathring{\Delta}_{d}}^{*}\phi^{j}_{i}=0$. Therefore the integral of $F^{*}(\omega^{i}_{j}\wedge\pi^{*}(\phi^{j}_{i}))$ over $\Phi^{-1}(S_{d}(T))=\Phi^{-1}(\bigcup\mathring{\Delta}_{d})$ is equal to zero.
@@
$\int_{\Phi^{-1}(S_{1}(T))}F^{*}(\omega^{i}_{j}\wedge\pi^{*}\phi^{j}_{i})=\int_{S_{1}(T)}f|_{S_{1}(T)}^{*}(\omega^{i}_{j}\wedge\pi^{*}(\phi^{j}_{i}))=\int_{S_{1}(T)}f|_{S_{1}(T)}^{*}\omega^{i}_{j}\wedge\phi^{j}_{i}.$
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle{f^{T}}^{*}(d\omega+\frac{1}{2}[\omega,\omega])\wedge\phi\right\rangle-\sum_{e\subset M}\int_{\mathring{e}}\left\langle[\![f^{T}|_{\mathring{e}}^{*}\omega]\!]\wedge\phi\right\rangle$ (10)
$\quad\quad-\sum_{p\subset M}\sum_{T\supset p}\int_{{\Phi^{T}}^{-1}(\mathring{p})}{F^{T}}^{*}\left\langle\omega^{i}_{j}\wedge\pi^{*}\phi^{j}_{i}\right\rangle.$
For the integrals over codimension-1 polytopes, let $\mathrm{I\!I}^{T}_{e}\in\Omega^{1}(e,\mathfrak{so}(n))$ be the form defined by $(\mathrm{I\!I}^{T}_{e})^{i}_{j}=0$ for all $1\leq i,j\leq n-1$ and $(\mathrm{I\!I}^{T}_{e})^{n}_{i}=\left\langle\nabla(E_{e}^{T})_{i},\vec{n}\right\rangle$ for $1\leq i\leq n$. We will use it to represent the second fundamental form of $e$. Also let $\tilde{\omega}_{e}^{T}=(E_{e}^{T})^{*}\omega$, that is, the full connection form for the frame $E_{e}^{T}$. Then the gauge transformation law (3) tells us that $f^{T}|_{e}^{*}\omega=\mathrm{Ad}((\mu_{e}^{T})^{-1})(\tilde{\omega}^{T}_{e})+(\mu_{e}^{T})^{-1}d\mu_{e}^{T}$.
@@
$=\mathrm{Ad}((\mu_{e}^{T})^{-1})(\tilde{\omega}^{T}_{e}-\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)(\tilde{\omega}^{T^{\prime}}_{e}))$
@@
Note also that here we have made key use of the fact that $i_{e}^{*}\phi=0$ for $e\subset\partial M$, so there is no issue with the boundary terms here. See section 4 for some discussion on how things change if $\phi$ is allowed to not pull back to zero on $\partial M$.

### v2

@@ 2.1 Conditions on Compatible Frames
An orthonormal frame $f=\bigsqcup_{T\subseteq M}f^{T}$ will be called compatible if it has some desirable properties that make $\left\langle\left\langle f^{*}\Omega_{\mathrm{dist}},\phi\right\rangle\right\rangle$ both correct from a geometrical standpoint and practical from a computational standpoint. The conditions that a compatible frame must satisfy are fairly technical and may seem arbitrary, so before stating them, we will first provide some motivation.
First, we want the individual vector fields in our frame to be “parallel” across codimension-1 polytopes, in some sense. For each interior codimension-1 polytope $\mathring{e}\subset\mathring{M}$, we can define a frame $E_{e}$ which is orthonormal in the metric $i_{e}^{*}g$; this metric is well-defined since $g$ has single-valued tangential-tangential components on $e$. If $e=T\cap T^{\prime}$, then $E_{e}$ can be extended to two orthonormal frames $E_{e}^{T}$ and $E_{e}^{T^{\prime}}$ by appending outward-facing normal vectors to $e$ for $g^{T}$ and $g^{T^{\prime}}$ respectively, which we denote by $\vec{n}$ and $\vec{n}^{\prime}$. Assume that for each $T$, $f^{T}$ can be continuously extended to $\mathring{e}$ by the $C^{1}$ section $f^{T}|_{\mathring{e}}:\mathring{e}\to\mathcal{F}_{O}(T)|_{\mathring{e}}$, and let $\mu_{e}^{T}:\mathring{e}\to O(n)$ be a map such that
@@
While somewhat arbitrary, this notion of “parallelism” is supported by the fact that a piecewise-smooth geodesic, defined as a locally energy-minimizing curve, must have a velocity vector that satisfies the same condition we have placed on the $f_{i}$’s.
@@
$=\sum_{i=1,2}\int_{\mathring{T_{i}}}\left(\left\langle{f^{T_{i}}}^{*}\omega\wedge d\phi\right\rangle+\frac{1}{2}\left\langle{f^{T_{i}}}^{*}[\omega,\omega]\wedge\phi\right\rangle\right)$
@@
As the Riemann curvature should be tensorial, it should always be true that changing the frame by $h$ results in an adjoint action by $h^{-1}$ on $f^{*}\Omega_{\mathrm{dist}}$. For this to be true for continuous $\phi$ and discontinuous $h$, it must be the case that $h^{-1}dh]\!]=0$ along every codimension-1 face $e=T\cap T^{\prime}$, so on $e$, $h^{T}=C_{e}h^{T^{\prime}}$ for some constant matrix $C_{e}$. We will restrict ourselves to the case $C_{e}=I$, since otherwise it would not be possible to modify the transformation $h$ so that it is the identity outside of a small neighborhood of $x_{0}$ but retains the jump condition, losing locality. What this means is that, if any particular frame $f$ is asserted to be compatible, then it is reasonable to assert that in a small enough neighborhood $U$ of a point $x\in\mathring{e}$, the set of compatible frames on $U$ (meaning restrictions of compatible frames to $U$) must be contained in $\{(f|_{U}\cdot h):h\;\text{is piecewise-smooth and continuous}\}$. Therefore, piecewise-smooth frames are usually not compatible, and the set of piecewise-smooth compatible frames must all have the same jump conditions along codimension-1 polytopes—the matrix $\mu_{e}^{T}(\mu_{e}^{T^{\prime}})^{-1}$ cannot depend on $f$. The choice $\mu_{e}^{T}(\mu_{e}^{T^{\prime}})^{-1}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}$ is the simplest, and it is consistent with the case of continuous metrics and frames.
@@
Note that since the final entries of $E_{e}^{T}$ and $E_{p,e}^{T}$ are associated to normal vectors to $e$ in the metric $g^{T}$, $A_{p,e}^{T}=\begin{bmatrix}A_{p,e}&0\\ 0&\pm 1\end{bmatrix}$, where $A_{p,e}$ does not depend on the polytope $T$. Because $E_{e}^{T^{\prime}}$ has a normal vector that points in a different direction to that of $E_{e}^{T}$, but $E_{p,e}^{T^{\prime}}$ does not, we also have $A_{p,e}^{T}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}A_{p,e}^{T^{\prime}}=A_{p,e}^{T^{\prime}}\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}$, and therefore
@@
which implies $\sum_{i=1}^{k}\theta_{p}^{T_{i}}=2m\pi$ for some integer $m$. However, in general, this sum can take any positive value if the metric is discontinuous at $p$. Therefore the frame $f^{T}$ cannot always be continuous at $p$. This is the origin of the angle defect. In order to control the discontinuity of $f^{T}$ as much as possible, we will restrict it to only rotate at a constant speed, and only in the plane orthogonal to $p$.
Summary of objects introduced.
The preceding paragraphs introduce some objects that are used throughout the rest of the paper. They are collected here for convenience.
Notation Definition
$E_{p}$ An arbitrary orthonormal frame on the codimension-2 polytope $p$. Entries are called $\tau_{1},\dots,\tau_{n-2}$.
$E_{e}$ An arbitrary orthonormal frame on the codimension-1 polytope $e$. Entries are called $E_{1},\dots,E_{n-1}$.
$\vec{n}^{T}$ The outward-pointing $g^{T}$-normal vector to a face $e\subset T$. Usually called just $\vec{n}$ when $T$ is implicitly known.
$\vec{\nu}$ The inward-pointing normal vector to a face $p\subset e$.
$E_{p,e}$ The $g^{e}$-orthonormal frame defined on $p$ with entries $(\tau_{1},\dots,\tau_{n-2},\vec{\nu})$
$A_{p,e}$ The map $p\to O(n-1)$ such that $E_{e}=E_{p,e}\cdot A_{p,e}$
$E_{e}^{T}$ The $g^{T}$-orthonormal frame defined on $e$ with entries $(E_{1},\dots,E_{n-1},\vec{n}^{T})$.
$E_{p,e}^{T}$ The $g^{T}$-orthonormal frame defined on $p$ with entries $(\tau_{1},\dots,\tau_{n-2},\vec{\nu},\pm\vec{n}^{T})$, where the sign on $\vec{n}^{T}$ is chosen so that $E_{p,e}^{T}$ is positively oriented in $M$.
$A_{p,e}^{T}$ The map $p\to O(n)$ such that $E_{e}^{T}=E_{p,e}^{T}\cdot A_{p,e}^{T}$.
$\mu_{e}^{T}$ The map $\mathring{e}\to O(n)$ such that $f^{T}|_{\mathring{e}}=E_{e}^{T}\cdot\mu_{e}^{T}$.
$\theta_{p}^{T}$ The map $p\to\mathbb{R}$ which measures the interior angle between the two faces $e,e^{\prime}\subset T$ whose intersection is equal to $p$.
To ensure that we are still able to apply integration by parts, despite the fact that the frame is discontinuous at $p$, we will follow the strategy outlined in [2, p. 2] and require that $f^{T}$ has some smoothness and continuity when pulled back to the blow-up $B_{T}$ of $T$. The blow-up $B_{T}$ of $T$ is essentially a polytope that has one codimension-1 face for each codimension-$d$ face of $T$ with $d\geq 1$, with tangencies related to inclusion relations between the original faces. There is a corresponding blow-down map $\Phi^{T}:B_{T}\to T$ which restricts to a diffeomorphism $\Phi^{T}|_{\mathring{B}_{T}}:\mathring{B}_{T}\to\mathring{T}$. The blow-up has an exceptional set which we will call $E_{B_{T}}$, consisting of the faces of $B_{T}$ of codimension $\geq 2$, and it can be safely ignored for the purpose of Riemann integration. Some information on blow-ups of manifolds can be found in the appendix.
@@
A frame $f=\bigsqcup_{T\subseteq M}f^{T}$, where $f^{T}:\mathring{T}\to\mathcal{F}_{O}(T)|_{\mathring{T}}$ is a $C^{2}$ orthonormal frame for the metric $g^{T}$, is compatible if:
@@
    For each $e\subset T\subseteq M$, let $\mu_{e}^{T}:\mathring{e}\to O(n)$ be the matrix such that $E_{e}^{T}\cdot\mu_{e}^{T}=f^{T}|_{\mathring{e}}$, and suppose $e=T\cap T^{\prime}$. Then $\mu_{e}^{T^{\prime}}=\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\mu_{e}^{T}$.
@@
    For each $p\subset T\subseteq M$, let $e\subset T$ be the unique face meeting $p$ such that the frame $E_{p,e}^{T}$ has an inward-pointing normal vector as its last entry. Then there exists an orientation-preserving embedding $\psi^{T}_{p}:[0,1]\times\mathring{p}\to\overline{{\Phi^{T}}^{-1}(\mathring{p})}$, where the orientation in $p$ is induced by the orientation on $e\subset T$, and the orientation on ${\Phi^{T}}^{-1}(\mathring{p})$ is induced from $B_{T}$, such that $\psi^{T}_{p}(\{0\}\times\mathring{p})\subset\overline{{\Phi^{T}}^{-1}(\mathring{e})}$ and $\Phi^{T}\circ\psi^{T}_{p}(s,x)=x$. Additionally, there exists a continuous function $r_{p,e}^{T}:p\to\mathbb{R}$ such that
@@
    for all $x\in\mathring{p}$ and all $s\in[0,1]$. In other words, the multi-valuedness of $f^{T}$ at codimension-2 faces is controlled so that $f^{T}$ is continuous when expressed in cylindrical coordinates around $x\in\mathring{p}$ and rotates in the plane orthogonal to $p$ at a rate depending only on $x$. There is no restriction on behavior near faces of higher codimension.
@@
As we mentioned above, condition 1 makes it possible to integrate by parts even though the frame may have discontinuities on polytopes of codimension 2 or greater. The fact that $F^{T}$ is Lipschitz and $C^{2}$ on the set $B_{T}\backslash E_{B_{T}}$ is necessary for the identity $d{F^{T}}^{*}\omega={F^{T}}^{*}d\omega$ to hold and for ${F^{T}}^{*}\omega$ to be bounded and continuously extendable to $B_{T}\backslash E_{B_{T}}$. These are all necessary conditions for Stokes’ Theorem to hold for the form ${F^{T}}^{*}(\omega\wedge\pi^{*}\phi)$ on $B_{T}$. Conditions 2 and 3 are necessary for the frame $f$ to have some semblance of continuity and for the distributional curvature to transform like a tensor, as mentioned previously, and condition 4 states that compatible frames can be “smoothed out” without introducing singularities.
@@
Having said all of this, one still might wonder if the use of $B_{T}$ could be avoided by performing integration by parts on $\overline{f^{T}(\mathring{T})}$ instead. After all, we are allowing $f^{T}$ to have discontinuities on codimension-2 faces of $T$, so it is reasonable to expect the boundary of $\overline{f^{T}(\mathring{T})}$ to have “extra” codimension-1 faces much like $B_{T}$ does. Unfortunately $\overline{f^{T}(\mathring{T})}$ may fail to be a Whitney manifold. For instance, this could happen if $r_{p,e}^{T}$ vanishes on a disconnected set which has nonzero measure in $p$. In a situation like this, integration by parts on $\overline{f^{T}(\mathring{T})}$ is infeasible. However, by pulling back forms by $F^{T}$ to the manifold $B_{T}$, which is regular enough to permit an integration by parts formula, we regain this ability even in the smooth category. Blow-ups are also a convenient tool for constructing compatible frames.
Remark 3.
We have ignored the question of whether the piecewise-smooth one-form $f^{*}\omega$ defined by ${f^{*}\omega}^{i}_{j}|_{\mathring{T}}=\left\langle\nabla^{T}f_{j}^{T},f_{i}^{T}\right\rangle_{T}$ on $T$ is “really” the connection form associated to the compatible frame $f$, despite the fact that $f$ is discontinuous. Here is one way to answer that question: Since the pullback of $f^{*}\theta^{i}$ to every element interface $e=T\cap T^{\prime}$ is single-valued, the distributional exterior derivative of $f^{*}\theta^{i}$ is simply its elementwise exterior derivative. Therefore the structure equations $df^{*}\theta^{i}=-f^{*}\omega^{i}_{j}\wedge f^{*}\theta^{j}$ hold in a distributional sense.
@@ 2.2 The Integration by Parts Step
For $d>2$, the integrals over $\Phi^{-1}(\mathring{\Delta}_{d})$ vanish for any codimension-$d$ face $\mathring{\Delta}_{d}\subset S_{d}(T)$. Indeed, $\phi^{j}_{i}$ is an (n2)(n-2)-form, so its trace $i_{\mathring{\Delta}_{d}}^{*}\phi^{j}_{i}$ vanishes on any $\mathring{\Delta}_{d}$ with $d>2$. Since $i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}F^{*}\pi^{*}=i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}\Phi^{*}=\Phi^{*}i_{\mathring{\Delta}_{d}}^{*}$, we have $i_{\Phi^{-1}(\mathring{\Delta}_{d})}^{*}F^{*}\pi^{*}\phi^{i}_{j}=\Phi^{*}i_{\mathring{\Delta}_{d}}^{*}\phi^{j}_{i}=0$. Therefore the integral of $F^{*}(\omega^{i}_{j}\wedge\pi^{*}\phi^{j}_{i})$ over $\Phi^{-1}(S_{d}(T))=\Phi^{-1}(\bigcup\mathring{\Delta}_{d})$ is equal to zero.
@@
$\int_{\Phi^{-1}(S_{1}(T))}F^{*}(\omega^{i}_{j}\wedge\pi^{*}\phi^{j}_{i})=\int_{S_{1}(T)}f|_{S_{1}(T)}^{*}(\omega^{i}_{j}\wedge\pi^{*}\phi^{j}_{i})=\int_{S_{1}(T)}f|_{S_{1}(T)}^{*}\omega^{i}_{j}\wedge\phi^{j}_{i}.$
@@
$=\sum_{T\subseteq M}\int_{\mathring{T}}\left\langle{f^{T}}^{*}(d\omega+\frac{1}{2}[\omega,\omega])\wedge\phi\right\rangle-\sum_{\mathring{e}\subset\mathring{M}}\int_{\mathring{e}}\left\langle[\![f^{T}|_{\mathring{e}}^{*}\omega]\!]\wedge\phi\right\rangle$ (10)
$\quad\quad-\sum_{\mathring{e}\subset\mathring{M}}\sum_{T\supset p}\int_{{\Phi^{T}}^{-1}(\mathring{p})}{F^{T}}^{*}\left\langle\omega\wedge\pi^{*}\phi\right\rangle.$
For the integrals over codimension-1 polytopes, let $\mathrm{I\!I}^{T}_{e}\in\Omega^{1}(e,\mathfrak{so}(n))$ be the form defined by $(\mathrm{I\!I}^{T}_{e})^{i}_{j}=0$ for all $1\leq i,j\leq n-1$ and $(\mathrm{I\!I}^{T}_{e})^{n}_{i}=\left\langle\nabla(E_{e}^{T})_{i},\vec{n}\right\rangle$ for $1\leq i\leq n$. We will use it to represent the second fundamental form of $e$. Also let $\tilde{\omega}_{e}^{T}=(E_{e}^{T})^{*}\omega$, that is, the full connection form for the frame $E_{e}^{T}$. Then the gauge transformation law (3) tells us that $f^{T}|_{\mathring{e}}^{*}\omega=\mathrm{Ad}((\mu_{e}^{T})^{-1})(\tilde{\omega}^{T}_{e})+(\mu_{e}^{T})^{-1}d\mu_{e}^{T}$.
@@
$=\mathrm{Ad}((\mu_{e}^{T})^{-1})\left(\tilde{\omega}^{T}_{e}-\mathrm{Ad}\left(\begin{bmatrix}I&0\\ 0&-1\end{bmatrix}\right)(\tilde{\omega}^{T^{\prime}}_{e})\right)$
@@
Note also that we have made key use of the fact that $i_{\mathring{e}}^{*}\phi=0$ for $e\subset\partial M$, so there is no issue with the boundary terms here. See Section 4 for some discussion on how things change if $\phi$ is allowed to not pull back to zero on $\partial M$.